a credit card bill for $680 was due on July 1st. Purchases of $315 were made on July 7th and $420 was charged on July 22nd. A payment of $400 was made on July 15th. find the average daily balance (in dollars) for the August 1st bill.
Answer #1

Balance in minus - = Payment made.

Balance in plus + = Expenses incurred

Jul 1-6 : credit card payment of $680 was due. (Bal 680$ )

Jul 7- 14 : Purchases were made of $ 315. (Bal : 680+ 315= 995$)

Jul 15-21 : Payment of $ 400. ( Bal : 995- 400 = 595$)

Jul 22-jul 31: Charges levied of $ 420. (Bal : 595+ 420 = 1015$)

Average Daily Balance as on Aug 1 : ( Closing Balance/ No of total days in billing cycle)

= $(680*6) + $(995*8)+$(595*7) + $(1015*11) /31

= $ (4080 + 7960 + 4165 + 11165)/31

= $ (27370)/31

= $ 882.90 ( opening average balance as on aug 1)